A student spent 60% of her money on books andspent 1/3 of the remainder on notebooks. She had $30 left. How much money did the student have at first?
Found 2 solutions by lwsshak3, ewatrrr:
Answer by lwsshak3(11628)

A student spent 60% of her money on books andspent 1/3 of the remainder on notebooks. She had $30 left. How much money did the student have at first?
***
let x=money student had at first.
60%x+40%x*(1/3)+30=x
.6x+.4x/3+30=x
lcd:3
1.8x+.4x+90=3x
2.2x+90=3x
.8x=90
x=112.50
money student had at first=$112.50

Answer by ewatrrr(24785)

T -.60T - .40T/3 = $30
40T - (1/3).40T = $30
(2/3).40T = $30
.40T = (3/2)30
.40T = 45
T = 45/.40 = $112.50
...
And...checking
.60T - .40T/3 + $30 = $67.50 + $15 + $30 = $112.50